# Nonlinear trident using WKB and worldline instantons

### v1

@@ II.6 Longitudinal limit
In this section we will consider a limit which is not usually considered. It was considered in [19] when considering backreaction on the field in Schwinger pair production, and in [20] when considering the trajectories of particles produced by Schwinger pair production in 4D e-dipole fields. In such cases, the electron can have a large momentum because it has been accelerated by an electric field. But, if the electron trajectory is parallel to the electric field and if the magnetic field vanishes along this trajectory, then regardless of how high the energy is, the electron will not see the field as a plane wave. Instead it will see a field on the form $E_{3}(t+z)$, which is very different from a plane wave even though they both depend on $t+z$. See [21, 22] for Schwinger pair production in such fields.
@@ III Space-time dependent fields with WKB
To determine the homogeneous solution $\delta S_{h}(\zeta[t,z])$, we consider the momentum
@@
where $b$ is another constant, which we can think of as an impact parameter. Comparing this with (155) shows that, if we evaluate the wave function $\phi(x^{\mu}[u])$ on the solution to the Lorentz-force equation, then $\zeta$ corresponds to $b$, so $\zeta$ would be approximately constant on a classical trajectory, and different values of $\zeta$ correspond to different impact parameters.
With $\zeta$ as a constant of motion, we find agreement between (160) and (162) if we identify the derivative of the homogeneous solution $\delta S_{h}^{\prime}(\zeta)$ with the constant $-c$ which appear in the momentum. To determine $c$, we argue that it is natural to choose it so that the exact momentum reduces to $p_{3}-A(t)-\delta A(t)$ when $\delta A$ does not depend on $z$. This implies $c=0$ and (up to an irrelevant constant) $\delta S_{h}=0$.
Thus, the correction to the exponential part of the electron wave function is given by (158) with $\delta S_{h}=0$,
$\delta S(t,z)=-\int_{0}^{t}\mathrm{d}t^{\prime}\frac{\pi_{3}(t^{\prime})}{\pi_{0}(t^{\prime})}\delta A_{3}\left(t^{\prime},z+\int_{t^{\prime}}^{t}\frac{\pi_{3}}{\pi_{0}}\right)\;.$ (164)
Comparing with the zeroth order (2), we see that the positron state is obtained by replacing $\pi_{3}=p_{3}-A\to-p_{3}-A$ and $\pi_{0}=\sqrt{m_{\scriptscriptstyle\perp}^{2}+(p_{3}-A)^{2}}\to-\sqrt{m_{\scriptscriptstyle\perp}^{2}+(p_{3}+A)^{2}}$.
If we take $t\to\pm\infty$ and $z$ to infinity with $\zeta$ kept finite, then
$\delta S(t,z)\to-\int_{0}^{\pm\infty}\mathrm{d}t^{\prime}\frac{\pi_{3}(t^{\prime})}{\pi_{0}(t^{\prime})}\delta A_{3}\left(t^{\prime},\zeta-\int_{0}^{t^{\prime}}\frac{\pi_{3}}{\pi_{0}}\right)\;,$ (165)
which is still a nontrivial function of $\zeta$. For a gauge field like (151), which always vanishes if $|z|\to\infty$ regardless of whether or not $\zeta$ is kept fixed, one might naively have expected that $S$ should not depend on $\delta A_{3}$, and in particular that $\delta S\to 0$. However, $\delta S$ is nonzero because of a memory effect, which we can see from the classical momentum (162). We already have a memory effect at zeroth order, since $\pi_{3}(-\infty)\neq\pi_{3}(+\infty)$ for $A_{3}(+\infty)\neq A_{3}(-\infty)$, so it is not strange that $\delta A_{3}(t,z)$ gives an additional change to the asymptotic momentum. But in order to find agreement between this classical prediction and $\partial_{\mu}\delta S$, the asymptotic $\delta S$ cannot vanish or be constant. While the asymptotic $\delta S$ has to solve (154) with $\delta A\to 0$, as we showed in (158), this still allows $\delta S$ to be a nontrivial function of $\zeta$. Moreover, it is not surprising that the classical asymptotic momentum depends nontrivially on the impact parameter, but, for such a nontrivial dependence to emerge from the wave function, and since $\zeta$ corresponds to the impact parameter, $\delta S$ must have a nontrivial dependence on $\zeta$.
@@ III.1 $\gamma_{z}\ll 1$ correction
$\delta A_{3}(t,z)=-(\gamma_{z}z)^{2}A_{3}(t)$ (167)
and from (164)
$\delta S=\gamma_{z}^{2}\int_{0}^{t}\mathrm{d}t^{\prime}\frac{\pi_{3}(t^{\prime})}{\pi_{0}(t^{\prime})}A_{3}(t^{\prime})\left(z+\int_{t^{\prime}}^{t}\frac{\pi_{3}}{\pi_{0}}\right)^{2}\;.$ (168)
@@
$I$ is given by (13) except that a factor of $1/E$ has been factored out, and $\varphi_{1}$ comes from $\delta S$. When we perform the time integrals with the saddle-point method, the saddle points will change due to $\varphi_{1}$, $t_{s}\approx t_{0}+\gamma_{z}^{2}t_{1}$. However, since $\varphi_{0}^{\prime}(t_{0})=0$, the correction $t_{1}$ drops out,
@@
Although $u$ and $v$ also appear in $\varphi_{1}$, we can perform these integrals in the same way as the time integrals, so the saddle point is simply $u=v=0$. If we had started with $\gamma_{z}=0$, then we would have found $\delta(\Sigma)$, so we expect that performing the $z$ integral for $\gamma_{z}>0$ will force $\Sigma$ to be small for $\gamma_{z}\ll 1$. Now the dependence on $z$ is still simple,
@@
$\exp\left\{\text{Re}\frac{(\gamma_{z}^{2}b-i\Sigma)^{2}}{2a\gamma_{z}^{2}}\right\}\;.$ (179)
We see from (30) and (33) that the zeroth order has a saddle point at $p_{3}=p_{3}^{\prime}=0$, i.e. at $\Sigma=0$. Since the saddle point for $t$ is imaginary, we see from (168) that $i\delta S(-p_{3})=[i\delta S(p_{3})]^{*}$, which means $a(-\Sigma)=[a(\Sigma)]^{*}$ and $b(-\Sigma)=[b(\Sigma)]^{*}$, so we can write the exponent in (179) as
$\frac{(\gamma_{z}^{2}b-i\Sigma)^{2}}{4a\gamma_{z}^{2}}+(\Sigma\to-\Sigma)\;.$ (180)
When performing the final-state momentum integrals, we therefore have a saddle point at $\Sigma=0$. Expanding around this point, we find, to leading order in $\gamma_{z}\ll 1$,
$\exp\left\{\frac{\gamma_{z}^{2}b^{2}}{2Ea}-\frac{\Sigma^{2}}{2Ea\gamma_{z}^{2}}\right\}\;,$ (181)
where $a$ and $b$ are now evaluated at $\Sigma=0$ and are therefore real. The width of an integrand around its saddle point scales in general as $\sqrt{E}$, but from the second term in (181) we see that the width for $\Sigma$ is even more narrow due to the $1/\gamma_{z}^{2}$ term. So if we take the limit $\gamma_{z}\to 0$ we have a regularized delta function which implies an approximate conservation of the longitudinal momentum, $\delta_{\rm reg}(p_{13}+p_{23}+p_{33}-p_{3})$.
Plugging $\Sigma=0$ into (178) gives a real saddle point for $z$. Note that it is nonzero and does not depend on $\gamma_{z}$. Since we consider symmetric fields, one might naively have expected the process to be centered in the middle of the field, i.e. with a saddle point at $z=0$. However, the initial state contains a single electron which is accelerated by the field in the $z$ direction, so the process is inherently non-symmetric. But it is only now when we consider $z$ dependent fields that we can fully see this non-symmetry.
Thus, the correction to the purely time-dependent result (30) is given by
$\exp\left\{-\frac{\gamma_{z}^{2}}{E}\left(c-\frac{b^{2}}{2a}\right)\right\}\;,$ (182)
where $a$, $b$ and $c$ are defined by (177). We will compare (178) and (182) with results obtained in the next section with the worldline instanton formalism, in which we do not have to expand in $\gamma_{z}$.
@@ IV Worldline instanton approach
We have developed an instanton approach based on open worldlines, which have allowed us to calculate the momentum spectrum of Schwinger pair production in space-time dependent fields [24, 20], nonlinear Compton scattering and Breit-Wheeler pair production [25, 26]. In all these cases there have been a single worldline. In the pair-production case we have a worldline that starts in the future as a positron, goes backwards in time, turns and goes to the future again as an electron. In the Compton case we have a worldline that starts in the past, as the incoming electron, and ends in the future as the outgoing electron. For trident we will have one worldline for the photon-emission step and another worldline for the pair-production step. Several steps are similar to the Breit-Wheeler case, so more details about those steps can be found in [26].
@@
As explained in [25], it is possible to choose different complex einbeins, i.e. different contours in the complex $u$ plane. One einbein that we found to be especially useful [20] starts (for the electron half of the B worldline) at $u=0$ and initially follows the imaginary axis until $u_{c}=-i|u_{c}|$, and afterwards goes parallel to the real axis towards $-i|u_{c}|+\infty$. To make the corner at $u_{c}$ smooth, we use the following einbein [24],
@@
where $L$ is approximately the length of the interval in which $\psi\approx 1$, and $W$ is the width of the of the regions where $\psi$ goes from 00 to $1$ or from $1$ to 00. $\epsilon=1$ for the final state particles, and $\epsilon=-1$ for the initial electron. We choose $W$ to be somewhat small, e.g. $W=0.1$. This is essentially just a choice of the shape of the bump function. $L$, on the other hand, is chosen such that the instanton becomes real asymptotically, so determining $L$ then requires a nontrivial computation. This choice of $L$ is not necessary. We would obtain the same probability and the same, real asymptotic momenta for other values of $L$, in particular for values which are held fixed even when we vary $\gamma$ or some other parameter, or even for different einbeins all together. But the asymptotic $q^{\mu}$ would not be real for general einbeins. This is not actually a problem, because we are only interested in the probability that the particles are produced with a certain (real) momenta, not that the particles appear at some position at some asymptotic time (though that could be an interesting problem for future studies). It can in fact be numerically convenient to choose such “non-physical” einbeins [24]. However, choosing $L$ such that $q^{\mu}$ is real asymptotically is not just conceptually appealing, it can also be useful for the calculations, e.g. to find approximations [20]. We will therefore choose such a “physical” einbein here, even though there is some extra numerical work to determine the values of $L$, which change when we change various parameters, such as $\gamma$.
$u=0$ corresponds to the middle of the instanton, where the emission or absorption happens. $u_{0}$ has a large and negative real part and corresponds to the incoming electron in the asymptotic past, or the outgoing positron in the asymptotic future. $u_{1}$ has a large and positive real part and corresponds to the outgoing electrons. The instantons are specified by asymptotic conditions at $u_{0}$ and $u_{1}$, which say that $z^{\prime}(\pm\infty)$ should be the momenta we measure and (for the physical einbein) $q^{\mu}$ should be real. In principle, we could find instantons for some arbitrary values of the momenta, but, since we do not have a good reason to choose otherwise, we will consider the saddle-point values of the final-state momenta. As shown in [24], the corresponding instantons have the following four conditions on the asymptotic velocities,
Im $\text{Im }z_{C,B}^{\prime}(u_{0,1})=0\;.$ (187)
To obtain the dependence on the initial momentum, we would replace Im $\text{Im }z_{C}^{\prime}(u_{0})=0$ with $z_{C}^{\prime}(u_{0})=-p_{3}$. The asymptotic energy, $t^{\prime}(u_{0,1})$, is automatically real since the instantons are on shell.
@@
Figure 1: Instanton quantities for $\gamma_{t}=2$ and $p_{1}=2$ as a function of the spatial inhomogeneity $\gamma_{z}$. $t_{C}[0;\gamma_{z}=0]=t_{B}[0;\gamma_{z}=0]$ is given by (196), so $t_{C}(u=0)=t_{B}(u=0)$ is purely imaginary. The momentum of the intermediate photon is also purely imaginary, Re $\text{Re }l_{3}=0$. $z[0;\gamma_{z}=0]$ is the WKB result (178) ($\Sigma=0$). In this particular case we have $a\approx 0.30$ and $b\approx 0.014$, so $z_{s}\approx-0.023$, so $z_{C}[u=0;\gamma_{z}]$ is purely real. From (200) and (201) we find $L_{C}^{\scriptscriptstyle-}(\gamma_{z}=0)\approx 0.25$ and $L_{C}^{\scriptscriptstyle+}(\gamma_{z}=0)=L_{B}^{\scriptscriptstyle\pm}(\gamma_{z}=0)\approx 0.51$. $\mathcal{A}$ is the exponential part of the probability, $P\propto e^{-\mathcal{A}/E}$, obtained from (202). $c_{1}$ is given by the WKB result (182).
@@
where $r$ is real. For moderately small $W$, the corner of the $u$ contour at $\sim u_{c}$ is smooth and $u$ does not actually become exactly equal to $u_{c}$ there. But Im $\text{Im }u$ does converge to Im $\text{Im }u_{c}$ for asymptotic $r$. For $|r|\gg 1$ we have
Im $\text{Im }u(r)=-\text{sign}(\epsilon r)L\coth\left[\frac{L}{W}\right]\;.$ (199)
@@
For sufficiently small $W$, the transcendental equation reduces to $L=-\text{Im }u_{c}$. But for $W=0.1$ and for the values of $\gamma$ and $p_{1}$ that we have considered, it is better to use the exact solutions of (200) and (201).

### v2

@@ II.6 Longitudinal limit
In this section we will consider a limit which is not usually considered. It was considered in [19] when considering backreaction on the field in Schwinger pair production, and in [23] when considering the trajectories of particles produced by Schwinger pair production in 4D e-dipole fields. In such cases, the electron can have a large momentum because it has been accelerated by an electric field. But, if the electron trajectory is parallel to the electric field and if the magnetic field vanishes along this trajectory, then regardless of how high the energy is, the electron will not see the field as a plane wave. Instead it will see a field of the form $E_{3}(t+z)$, which is very different from a plane wave even though they both depend on $t+z$. See [24, 25] for Schwinger pair production in such fields.
@@ III Space-time dependent fields with WKB
To interpret the homogeneous solution $\delta S_{h}(\zeta[t,z])$, we consider the momentum
@@
where $b$ is another constant, which we can think of as an impact parameter. Comparing this with (155) shows that, if we evaluate the wave function $\phi(x^{\mu}[u])$ on the solution to the Lorentz-force equation, then $\zeta$ corresponds to $b$, so $\zeta$ would be approximately constant on a classical trajectory, and different values of $\zeta$ correspond to different impact parameters. With $\zeta$ as a constant of motion, we find agreement between (160) and (162) if we identify the derivative of the homogeneous solution $\delta S_{h}^{\prime}(\zeta)$ with the constant c-c which appears in the momentum.
We determine $\delta S_{h}^{\prime}(\zeta)$ by demanding that the zeroth order already gives the correct asymptotic momentum, which gives
$\delta S(t,z)=-\int_{t_{a}}^{t}\mathrm{d}t^{\prime}\frac{\pi_{3}(t^{\prime})}{\pi_{0}(t^{\prime})}\delta A_{3}\left(t^{\prime},z+\int_{t^{\prime}}^{t}\frac{\pi_{3}}{\pi_{0}}\right)\;,$ (164)
where $t_{a}\to\infty$ ($t_{a}\to$-$\infty$) for an outgoing (incoming) electron. By comparing this with the zeroth order (2), we see that the positron state is obtained by replacing $\pi_{3}=p_{3}-A\to-p_{3}-A$ and $\pi_{0}=\sqrt{m_{\scriptscriptstyle\perp}^{2}+(p_{3}-A)^{2}}\to-\sqrt{m_{\scriptscriptstyle\perp}^{2}+(p_{3}+A)^{2}}$.
If we take $t\to-t_{a}$ and $z$ to infinity with $\zeta$ kept finite, then
$\delta S(t,z)\to-\int_{t_{a}}^{-t_{a}}\mathrm{d}t^{\prime}\frac{\pi_{3}(t^{\prime})}{\pi_{0}(t^{\prime})}\delta A_{3}\left(t^{\prime},\zeta-\int_{0}^{t^{\prime}}\frac{\pi_{3}}{\pi_{0}}\right)\;,$ (165)
which is still a nontrivial function of $\zeta$. For a gauge field such as (151), which always vanishes if $|z|\to\infty$ regardless of whether $\zeta$ is kept fixed, one might naively have expected that $\delta S\to 0$ in this limit. However, $\delta S$ is nonzero because of a memory effect, which we can see from the classical momentum (162). We already have a memory effect at zeroth order, since $\pi_{3}(-\infty)\neq\pi_{3}(+\infty)$ for $A_{3}(+\infty)\neq A_{3}(-\infty)$, so it is not strange that $\delta A_{3}(t,z)$ gives an additional change to the asymptotic momentum. But in order to find agreement between this classical prediction and $\partial_{\mu}\delta S$, the asymptotic $\delta S$ cannot vanish or be constant. While the asymptotic $\delta S$ has to solve (154) with $\delta A\to 0$, as we showed in (158), this still allows $\delta S$ to be a nontrivial function of $\zeta$. Moreover, it is not surprising that the classical asymptotic momentum depends nontrivially on the impact parameter, but, for such a nontrivial dependence to emerge from the wave function, and since $\zeta$ corresponds to the impact parameter, $\delta S$ must have a nontrivial dependence on $\zeta$.
@@ III.1 $\gamma_{z}\ll 1$ correction
$\delta A_{3}(t,z)=-(\gamma_{z}z)^{2}A_{3}(t)\;.$ (167)
However, after making this expansion, we no longer have $\delta A_{3}(t,z)\to 0$ as $|z|\to\infty$. So, before we expand, we perform a partial integration in (164),
$\delta S(t,z)=-\int_{t_{a}}^{t}\mathrm{d}t^{\prime}E(t^{\prime})f\left(z+\int_{t^{\prime}}^{t}\frac{\pi_{3}}{\pi_{0}}\right)\;,$ (168)
where $E(t)=A_{3}^{\prime}(t)$ and $f^{\prime}(z)=F(z)-1$. This is also what we would have obtained if we had started in a gauge with $\delta A_{0}\neq 0$ rather than $\delta A_{3}\neq 0$. Now we can expand,
$\delta S=\frac{\gamma_{z}^{2}}{3}\int_{t_{a}}^{t}\mathrm{d}t^{\prime}E(t^{\prime})\left(z+\int_{t^{\prime}}^{t}\frac{\pi_{3}}{\pi_{0}}\right)^{3}\;.$ (169)
@@
II is given by (13) except that a factor of $1/E$ has been factored out, and $\varphi_{1}$ comes from $\delta S$. When we perform the time integrals with the saddle-point method, the saddle points will change due to $\varphi_{1}$, $t_{s}\approx t_{0}+\gamma_{z}^{2}t_{1}$. However, since $\varphi_{0}^{\prime}(t_{0})=0$, the correction $t_{1}$ drops out,
@@
Although $u$ and $v$ also appear in $\varphi_{1}$, we can perform these integrals in the same way as the time integrals, so the saddle point is simply $u=v=0$.
If we had started with $\gamma_{z}=0$, then we would have found $\delta(\Sigma)$, so we expect that performing the $z$ integral for $\gamma_{z}>0$ will force $\Sigma$ to be small for $\gamma_{z}\ll 1$. Now the dependence on $z$ is still simple,
@@
$\exp\left\{-\frac{\gamma_{z}^{2}}{E}\left(2c_{r}-\frac{b_{r}^{2}}{2a_{r}}+\frac{a_{r}}{2|a|^{2}}\left[\frac{\Sigma}{\gamma_{z}^{2}}-\frac{a_{r}b_{i}-a_{i}b_{r}}{a_{r}}\right]^{2}\right)\right\}\;,$ (180)
where $a_{r}=\operatorname{Re}a$ and $a_{i}=\operatorname{Im}a$ etc.
As $\gamma_{z}\to 0$, (180) gives a regularized delta function which implies an approximate conservation of the longitudinal momentum, $\delta_{\rm reg}(p_{13}+p_{23}+p_{33}-p_{3})$. We also see that for $\gamma_{z}\neq 0$, the saddle point is nonzero,
$\Sigma_{s}=\gamma_{z}^{2}\frac{a_{r}b_{i}-a_{i}b_{r}}{a_{r}}\;.$ (181)
Plugging (181) into (179) gives a real saddle point for $z_{s}$,
$z_{s}(\Sigma_{s})=-\frac{b_{r}}{2a_{r}}\;.$ (182)
Note that $z_{s}$ is nonzero and does not depend on $\gamma_{z}$. Since we consider symmetric fields, one might naively have expected the process to be centered in the middle of the field, i.e. with a saddle point at $z_{s}=0$. However, the initial state contains a single electron which is accelerated by the field in the $z$ direction, so the process is inherently non-symmetric. But it is only now when we consider $z$ dependent fields that we can fully see this non-symmetry.
For $\gamma_{z}=0$ we have a saddle point at ${\bf p}_{1}={\bf p}_{2}={\bf p}_{3}={\bf p}/3$. But for $\gamma_{z}>0$, this symmetry is broken. For the longitudinal component, we see this from (181). For the $x$ component, it follows because if we expand (180) around what would have been a saddle point for $\gamma_{z}=0$, then we find a linear term,
$\exp\left\{-\frac{Y}{\omega}\delta P_{1}^{2}+\frac{2\gamma_{z}^{2}}{E}\frac{p_{11}}{m_{1}}\frac{b_{r}}{a_{r}}\frac{\partial b_{r}}{\partial m_{1}}\delta P_{1}\right\}\;,$ (183)
where $Y$ is the coefficient of $\delta P_{1}^{2}$ in the first row of (31). This linear term implies that the saddle point is shifted from $\delta P_{1}^{s}=0$ to $\delta P_{1}^{s}\neq 0$. Since there is no linear term in $\delta p_{1}$, we still have $\delta p_{1}^{s}=0$, so the saddle points for the two final-state electrons are still equal, $p_{11}^{s}=p_{21}^{s}$, but they are no longer equal to the positron momentum, $p_{11}^{s}=p_{21}^{s}\neq p_{31}^{s}$. We find
$p_{11}^{s}=p_{21}^{s}=\frac{p_{1}}{3}+\delta P_{1}^{s}\qquad p_{31}^{s}=\frac{p_{1}}{3}-\delta P_{1}^{s}\;,$ (184)
where $p^{\prime}_{1}=p_{1}/3$ is the saddle point to zeroth order, and
$\delta P_{1}^{s}=\frac{p^{\prime}_{1}}{m^{\prime}}\frac{b_{r}}{a_{r}}\frac{\partial b_{r}}{\partial m^{\prime}}\frac{\gamma_{t}\gamma_{z}^{2}}{Y}\;.$ (185)
We will compare the $\mathcal{O}(\gamma_{z}^{2})$ results in this section with the results obtained in the next section with the worldline instanton formalism, in which we do not have to expand in $\gamma_{z}$.
@@ IV Worldline instanton approach
We have developed an instanton approach based on open worldlines, which has allowed us to calculate the momentum spectrum of Schwinger pair production in space-time dependent fields [27, 23], nonlinear Compton scattering and Breit-Wheeler pair production [28, 29]. In all these cases there has been a single worldline. In the pair-production case we have a worldline that starts in the future as a positron, goes backwards in time, turns and goes to the future again as an electron. In the Compton case we have a worldline that starts in the past, as the incoming electron, and ends in the future as the outgoing electron. For trident we will have one worldline for the photon-emission step and another worldline for the pair-production step. Several steps are similar to the Breit-Wheeler case, so more details about those steps can be found in [29].
@@
As explained in [28], it is possible to choose different complex einbeins, i.e. different contours in the complex $u$ plane. One einbein that we found to be especially useful [23] starts (for the electron half of the B worldline) at $u=0$ and initially follows the imaginary axis until $u_{c}=-i|u_{c}|$, and afterwards goes parallel to the real axis towards $i|u_{c}|+\infty$. To make the corner at $u_{c}$ smooth, we use the following einbein [27],
@@
where $L$ is approximately the length of the interval in which $\psi\approx 1$, and $W$ is the width of the regions where $\psi$ goes from 0 to $1$ or from $1$ to 0. $\epsilon=1$ for the final state particles, and $\epsilon=-1$ for the initial electron. We choose $W$ to be somewhat small, e.g. $W=0.1$. This is essentially just a choice of the shape of the bump function. $L$, on the other hand, is chosen such that the instanton becomes real asymptotically, so determining $L$ then requires a nontrivial computation. This choice of $L$ is not necessary. We would obtain the same probability and the same, real asymptotic momenta for other values of $L$, in particular for values which are held fixed even when we vary $\gamma$ or some other parameter, or even for different einbeins all together. But the asymptotic $q^{\mu}$ would not be real for general einbeins. This is not actually a problem, because we are only interested in the probability that the particles are produced with a certain (real) momenta, not that the particles appear at some position at some asymptotic time (though that could be an interesting problem for future studies). It can in fact be numerically convenient to choose such “non-physical” einbeins [27]. However, choosing $L$ such that $q^{\mu}$ is real asymptotically is not just conceptually appealing, it can also be useful for the calculations, e.g. to find approximations [23]. We will therefore choose such a “physical” einbein here, even though there is some extra numerical work to determine the values of $L$, which change when we change various parameters, such as $\gamma$.
$u=0$ corresponds to the middle of the instanton, where the emission or absorption happens. $u_{0}$ has a large and negative real part and corresponds to the incoming electron in the asymptotic past, or the outgoing positron in the asymptotic future. $u_{1}$ has a large and positive real part and corresponds to the outgoing electrons. The instantons are specified by asymptotic conditions at $u_{0}$ and $u_{1}$, which say that $z^{\prime}(\pm\infty)$ should be the momenta we measure and (for the physical einbein) $q^{\mu}$ should be real. We can find instantons for some arbitrary values of the momenta by imposing the asymptotic conditions,
$z_{C}^{\prime}(\infty)=-p_{13}\qquad z_{B}^{\prime}(\infty)=-p_{23}\qquad z_{B}^{\prime}(-\infty)=p_{33}\;.$ (190)
However, since we do not have a good reason to choose otherwise, we will consider the saddle-point values of the final-state momenta, and calculate the Hessian matrix to determine the widths of the spectrum around the saddle point. As shown in [27], the corresponding instantons have the following four conditions on the asymptotic velocities,
$\text{Im }z_{C,B}^{\prime}(u_{0,1})=0\;.$ (191)
To obtain the dependence on the initial momentum, we would replace $\text{Im }z_{C}^{\prime}(u_{0})=0$ with $z_{C}^{\prime}(u_{0})=-p_{3}$. The asymptotic energy, $t^{\prime}(u_{0,1})$, is automatically real since the instantons are on shell.
@@
Figure 1: Instanton quantities for $\gamma_{t}=2$ and $p_{1}=2$ as a function of the spatial inhomogeneity $\gamma_{z}$. $t_{C}[0;\gamma_{z}=0]=t_{B}[0;\gamma_{z}=0]$ is given by (200), so $t_{C}(u=0)=t_{B}(u=0)$ is purely imaginary. The momentum of the intermediate photon is also purely imaginary, $\text{Re }l_{3}=0$. $z[0;\gamma_{z}=0]$ is the WKB result (179) ($\Sigma=0$). In this particular case we have $a\approx 0.30$ and $b\approx 0.014$, so $z_{s}\approx-0.023$, so $z_{C}[u=0;\gamma_{z}]$ is purely real. From (204) and (205) we find $L_{C}^{\scriptscriptstyle-}(\gamma_{z}=0)\approx 0.25$ and $L_{C}^{\scriptscriptstyle+}(\gamma_{z}=0)=L_{B}^{\scriptscriptstyle\pm}(\gamma_{z}=0)\approx 0.51$. $\mathcal{A}$ is the exponential part of the probability, $P\propto e^{-\mathcal{A}/E}$, obtained from (206). $c_{1}$ is given by the WKB result (180).
@@
where $r$ is real. For moderately small $W$, the corner of the $u$ contour at $\sim u_{c}$ is smooth and $u$ does not actually become exactly equal to $u_{c}$ there. But $\text{Im }u$ does converge to $\text{Im }u_{c}$ for asymptotic $r$. For $|r|\gg 1$ we have
$\text{Im }u(r)=-\text{sign}(\epsilon r)L\coth\left[\frac{L}{W}\right]\;.$ (203)
@@
[…]